# Visualization Tasks for Unlabelled Graphs

### v1

@@ Abstract
We investigate tasks that can be accomplished with unlabelled graphs, where nodes do not have persistent or semantically meaningful labels. New techniques to visualize these graphs have been proposed, but more understanding of unlabelled graph tasks is required before they can be adequately evaluated. Some tasks apply to both labelled and unlabelled graphs, but many do not translate between these contexts. We propose a taxonomy of unlabelled graph abstract tasks, organized according to the Scope of the data at play, the Action intended by the user, and the Target data under consideration. We show the descriptive power of this task abstraction by connecting to concrete examples from previous frameworks, and connect these abstractions to real-world problems. To showcase the evaluative power of the taxonomy, we perform a preliminary assessment of 6 visualizations for each task. For each combination of task and visual encoding, we consider the effort required from viewers, the likelihood of task success, and how both factors vary between small-scale and large-scale graphs. Availability: Our supplemental materials are available at osf.io/e23mr
keywords:
Task taxonomy, network structure
Introduction
@@
Task taxonomies support the analysis, design, and evaluation of visualization tools [12]. In the case of graph-structured data, where nodes are interconnected through edges, tasks may pertain to the topological structure of that connectivity, or attributes attached to nodes or edges [35]. In nearly all previous visualization contexts, graphs are assumed to be labelled graphs: that is, they are assumed to have persistent unique identifiers for the nodes, even when the graph has no other attributes.
@@
However, there are network visualization approaches that do not make this assumption, and rely purely on topological information that is completely independent of node labels. These unlabelled graph methods emphasize properties of graph data structures that remain unchanged when node labels are ignored, reassigned, or shuffled. These approaches are sometimes called invariant descriptor idioms [47]; examples include BMatrix Network Portraits [6], Graph Thumbnails [70], and Census Plots [47].
A major challenge to evaluating these approaches is that no previous task taxonomy for networks is oriented towards the unlabelled graph perspective. Enumerating and categorizing what tasks can be performed on unlabelled graph data is an open problem that needs to be addressed before human-subjects studies can be run. While some tasks apply to both labelled and unlabelled graphs, many tasks designed for the labelled context do not translate to graphs under the unlabelled constraint. Conversely, some tasks that are challenging with traditional graph visualization approaches become feasible with visual encodings of unlabelled graphs.
@@
We attack this problem through a combination of literature review and reflective synthesis. We identify which of the many concrete tasks discussed in previous work pertain to unlabelled graphs, considering both abstracted tasks stated explicitly in the visualization literature and implicit tasks that we infer from theoretical works, practitioner surveys, and domain-specific literature. To make these choices, we create a data abstraction that allows us to quickly exclude many tasks according to stated dataset characteristics.
We propose a three-part task abstraction that maps concrete unlabelled graph tasks into a taxonomy of Scope + Action + Target triplets. The Scope ranges from multiple graphs through single graphs down to constituent elements, and determines the available set of targets at each level. The Target is a noun identifying which aspects of the data are relevant, such as scale-free network structure, or the dense interconnections of a clique, or the local topological neighborhood of a hub node. The Action is a verb that defines a user’s goal, such as compare, classify, or quantify.
The primary contribution of this paper is our Scope + Action + Target taxonomy of abstract visualization tasks for unlabelled graphs (Sec. 3). The data abstraction that shaped this task taxonomy is a secondary contribution (Sec. 2).
@@
After framing our work with respect to the existing literature (Sec. 5), we validate our task taxonomy through consideration of its descriptive, generative, and evaluative power [8]. We demonstrate its descriptive power by mapping previously proposed concrete tasks into these abstract triplets. We also provide some preliminary evidence of the ecological validity of the concrete tasks by connecting them to real-world problems stated in domain literature. Conversely, we show the generative power of our taxonomy by proposing new concrete tasks from the combinatorial possibilities of these triplets. To showcase the evaluative power of our taxonomy, we perform a preliminary assessment of 6 visual encodings, both traditional and invariant descriptor views, in relation to 17 concrete tasks (Sec. 6). In this exercise, we consider the effort demanded from viewers and the likelihood of success for both small and large graphs.
1 Process
We interleaved five activities in parallel, through multiple rounds of iterative refinement that occurred across six months, through individual work from the first author and weekly meetings with two other authors, until convergence where consensus was reached. Our process included the methods of literature review and reflective synthesis.
First Author Title Tasks Tax Year Ref
@@ 1 Process
Figure 1: We report on the 64 papers in our literature review, providing first author name and titles for each. The list is organized according to whether we chose previously articulated tasks, gleaned implicitly mentioned tasks, or excluded all tasks from the paper (Tasks). We document whether the paper features a task taxonomy (Tax), paper Year, and provide the citation (Ref).
Figure 2: Data abstraction model with an illustrative example. (A) In our data abstraction model, nested layers progressively decrease in semantic richness, shown here from high on the left to low on the right: Formalism is a stand-in term for named graph models with rich associated semantics, Attribute allows for non-topological semantic information to be associated with nodes or edges, Label indicates that there are persistent unique identifiers that can convey semantic meaning, and Core is the innermost layer where all semantic information is stripped away, leaving only unlabelled graphs with the connectivity relationships of the core graph topology. (B) As an illustrative example, drawn in red are the 6 unlabelled graphs of 4 nodes. If persistent labels (here shown as node colors) exist, as in the Label layer, this set of 6 graphs expands to the set of 38 labelled graphs of 4 nodes. Within each of the 6 groupings based on graph topology (red enclosures) the graphs are structurally equivalent, and distinguishable from each other only through the meaning conveyed through the node labels.
1.1 Curating Concrete Tasks
To identify concrete tasks from previous work, the first author conducted a literature review of 64 publications that touched on networks and/or visualization, using a combination of network visualization task taxonomies and surveys already known to three of the authors for seed papers, and broad searches for potentially relevant material. Of these papers, 40 yielded relevant tasks. Of these, 14 papers included explicit abstract tasks that we could simply extract. For example, “Is the node an articulation node?” is listed in Lee et al’s seminal taxonomy of network visualization tasks [35].
From 26 other papers we gleaned end-user tasks, either inferring them through definitions of graph-theoretic concepts and terminology or descriptions of the use cases for algorithms or techniques. For example, Newman’s landmark 2003 paper on complex network structure [44] implicitly suggests that “finding the length of the longest geodesic” and “counting the number of longest geodesics” are relevant tasks to draw insights about the network’s maximum traversable distances, and the induced paths that support these traversals.
Fig. 1 summarizes the results, coded by whether tasks were chosen, gleaned, or excluded. We also provide a spreadsheet version as supplemental material.
@@ 1.2 Developing Data Abstraction
As we considered which concrete tasks to include and which to exclude, we developed a data abstraction through reflective synthesis that could serve as a filter to exclude tasks that cannot be undertaken with unlabelled graphs. This data abstraction allowed us to exclude 24 papers based on data characteristics (Fig. 1). We also used our data abstraction to identify which aspects of specific tasks should be extracted.
@@ 1.3 Mapping Task Abstraction
We developed the task abstraction that covered the full suite of relevant concrete tasks with a concise set of terms through multiple rounds of reflective synthesis. We finalized the framework of three levels (Scope, Action, Target) early on, but refined the set of items at each level extensively.
@@ 1.4 Analyzing Visual Encodings
We considered the strengths and weaknesses of different visual encodings for specific concrete tasks using reflective synthesis, considering the encoding results across an eclectic collection of 81 networks [47], drawing on the experience and expertise of the authors. Two are doctoral students, and two are senior visualization researchers.
@@ 1.5 Generating Concrete Tasks
We generated new concrete tasks for abstraction triplets where we had not found compelling examples articulated in prior work, again using reflective synthesis.
@@ 2 Data Abstraction
Formalism. This outermost layer of our data abstraction (Fig. 2-A) addresses the rich semantic complexity of real-world networked data. We define Formalism as a stand-in for named graph models, which have specific facets of associated semantic data that help us interpret and analyze the model. For example, dynamic graphs have node and edge attributes parsed as time slices [1, 30, 9], compound graphs can have disjoint or overlapping groups assigned to nodes [57, 65], hypergraphs require a generalized definition of relationship that allows more than two nodes to be connected by the same edge at once [20], and heterogeneous graphs have multiple different kinds of attributes associated to nodes and edges [52, 45, 14].
@@
Formalisms cannot be unlabelled graphs by definition: they explicitly associate semantic information beyond purely topological structure with network data. Any task that pertains to the use of those formalism semantics can easily be ruled out as impossible to achieve with an unlabelled graph.
@@
Attribute. Even graphs without the semantic complexity of formalisms can contain Attributes assigned to nodes and edges, which encode domain-specific information that augments the topological structure. For example, in a social network, semantic data can include node attributes such as a person’s height and age [44], while in a biological system such semantic data may take the form of edge attributes representing the reactivity of different metabolites [43]. Attribute data is used extensively in graph drawing [32].
@@
These attributes on nodes or edges are independent from and orthogonal to the topological structure of the network. Signed values (positive or negative) and scalar weights (for nodes and/or edges) are special cases of attributes. Any graph that falls into this layer cannot be an unlabelled graph, because any task that relies on the use of attributes can be ruled out as impossible to achieve with an unlabelled graph.
@@
Label. The next lowest layer of semantic richness is graphs with Labels, which we define as semantically meaningful unique identifiers that are persistent and can serve to distinguish nodes from each other. Although this particular distinction is not common, we propose it because even graphs that are typically considered as being attribute-free in the visualization literature do in fact have labels.
Although labels could be considered as a special case of attributes, we distinguish them from the more general attributes in the previous layer because they are at the heart of the distinction between labelled and unlabelled graphs. Any task that relies on given labels is inapplicable to unlabelled graphs.
@@
Directed graphs require labels to establish persistent edge direction, which conversely implies that unlabelled graphs cannot be directed. Similarly, unlabelled graphs cannot have self loops.
@@
Core. At the innermost layer, outlined in red in Fig. 2-A, is the connectivity information of networked data, which establishes how nodes are connected via edges, or the Core graph topology. In this layer, nodes can be distinguished from each other only in terms of topological connectivity, not through persistent names, so there is no semantic information at all at this layer.
@@
To illustrate, Fig. 2-B compares labelled and unlabelled graphs: with 4 nodes there are only 6 different unlabelled graphs (node-link views in red), while there are 38 possible labelled graphs, which we show with node colors to indicate the labels, rather than the more typical approach of using text strings (which would be too small to read). The correspondence with the unlabelled graphs is shown through the 6 groupings (red enclosures) in Fig. 2-B, each of which has a respective unlabelled graph. Although the topological structure is the same within the groups, the labelled graphs are distinguishable from each other through the persistent labels that can convey meaning. For example, in the third red enclosure from the right, with 4 labelled graphs, the single degree-3 node can be bound to each of the different labelled cases, shown with the different colors. Without labels there is no difference between the four graphs within that group, so we have only a single unlabelled graph corresponding to each grouping. In more mathematical terms, the unlabelled graph is the isomorphism class of the labelled graph [3].
@@
Although we use the term unlabelled graph throughout this paper, following the conventions of previous work, it is arguably a misnomer in practice: to instantiate a graph in any computational context, such as to load data or perform a hop-based traversal of it, there must be some kind of unique identifiers that would serve as labels. However, within the Core layer, these identifiers are non-persistent and arbitrary. To clarify, what we emphasize in this paper with our data abstraction model is not the presence of labels, but whether the labels carry semantic meaning.

### v2

@@ Abstract
We investigate tasks that can be accomplished with unlabelled graphs, which are graphs with nodes that do not have attached persistent or semantically meaningful labels. New visualization techniques to represent unlabelled graphs have been proposed, but more understanding of unlabelled graph tasks is required before these techniques can be adequately evaluated. Some tasks apply to both labelled and unlabelled graphs, but many do not translate between these contexts. We propose a data abstraction model that distinguishes the Unlabelled context from the increasingly semantically rich Labelled, Attributed, and Augmented contexts. We filter tasks collected and gleaned from the literature according to our data abstraction and analyze the surfaced tasks, leading to a taxonomy of abstract tasks for unlabelled graphs. Our task taxonomy is organized according to the Scope of the data at play, the Action intended by the user, and the Target data under consideration. We show the descriptive power of this task abstraction by connecting to concrete examples from previous frameworks, and connect these abstractions to real-world problems. To showcase the evaluative power of the taxonomy, we perform a preliminary assessment of 6 visualizations for each task. For each combination of task and visual encoding, we consider the effort required from viewers, the likelihood of task success, and how both factors vary between small-scale and large-scale graphs. Availability: Our Supplemental materials are available at osf.io/e23mr.
Index Terms:
Task taxonomy, unlabelled graphs, invariant plots
1 Introduction
@@
Task taxonomies support the analysis, design, and evaluation of visualization tools [1]. In the case of graph-structured data, where nodes are interconnected through edges, tasks may pertain to the topological structure of that connectivity, or attributes attached to nodes or edges [2]. In nearly all previous visualization contexts, graphs are assumed to be labelled graphs; that is, they have persistent unique identifiers for the nodes, even when the graph has no other attributes.
@@
However, there are network visualization approaches that do not make this assumption, and rely purely on topological information that is completely independent of node labels. These unlabelled graph methods emphasize properties of graph data structures that remain unchanged when node labels are ignored, reassigned, or shuffled. These techniques have been called invariant descriptor visualization idioms [3], which we shorten here to invariant plots. Examples include the BMatrix Network Portrait [4], Graph Thumbnails [5], and Census plots [3].
Researchers may find it difficult to persuasively evaluate the efficacy and utility of invariant plots until a task taxonomy that pertains to them is introduced. Enumerating and categorizing what tasks can be performed on unlabelled graph data is an open problem that needs to be addressed before human-subjects studies can be run. The type of tasks supported by these invariant plots are fundamentally different from those targeted by the traditional views dedicated to semantically meaningful and persistent labels: while some tasks apply to both labelled and unlabelled graphs, many tasks designed for the labelled context do not translate to the unlabelled context. For example, the task of following the spread of disease between people through a contagion network presupposes persistent node labels to identify those specific people [6]. Conversely, some tasks that are challenging with traditional graph visualization approaches become feasible with visual encodings of unlabelled graphs. For example, pairwise comparison between two graphs could be easier with invariant plots that exploit a shared absolute coordinate system, than with node-link layouts where the spatial position of node points is arbitrary and unconstrained [3].
@@
We attack this problem through a combination of literature review and reflective synthesis (Sec. 2). We identify which of the many concrete tasks discussed in previous work pertain to unlabelled graphs, considering both abstract tasks stated explicitly in the visualization literature and implicit tasks that we infer from theoretical works, practitioner surveys, and domain-specific literature. To make these choices, we create a data abstraction that assigns graphs into the categories of Augmented, Attributed, Labelled, and Unlabelled (Sec. 3), which allows us to filter tasks from existing literature (Sec. 4). We then propose our task taxonomy, which is an abstraction that maps concrete unlabelled graph tasks into Scope + Action + Target triplets (Sec. 5).
@@
We show the generative power of our taxonomy by proposing new concrete tasks from the combinatorial possibilities of these triplets (Sec. 6.2 and 6.3). We also validate our task taxonomy through consideration of its descriptive, generative, and evaluative power [7]. We demonstrate its descriptive power by mapping many previously proposed concrete tasks into these abstract triplets; and provide some preliminary evidence of the ecological validity of the concrete tasks by connecting them to real-world problems stated in domain literature (Sec. 6.4). To showcase the descriptive and evaluative power of our taxonomy, we provide an extended example of its use in a preliminary assessment of 6 visual encodings, both traditional and invariant descriptor views, in relation to 17 concrete tasks (Sec. 7). In this exercise, we made educated guesses about the effort demanded from viewers and the likelihood of success for both small and large graphs. This assessment is not intended to provide final answers to these questions, but rather to showcase the utility of the taxonomy for the design of future user studies.
In summary, the primary contribution of this paper is our Scope + Action + Target taxonomy of abstract visualization tasks for unlabelled graphs (Sec. 5); the data abstraction model behind the construction of this task taxonomy is a secondary contribution (Sec. 3).
First Author Title Tasks Taxonomy Year Ref
@@
Pandey A state-of-the-art survey of tasks for tree design and evaluation with a curated task dataset Choose X 2022 [12]
@@
Figure 1: We report on the 65 papers in our literature review, providing first author name and titles for each. The list is organized according to whether we chose previously articulated tasks, gleaned implicitly mentioned tasks, or excluded all tasks from the paper (Tasks). We document whether the paper features a task taxonomy (Taxonomy), publication (Year), and provide the citation (Ref).
2 Process
We interleaved six process stages through multiple rounds of iterative refinement that occurred across six months, through individual work from the first author and weekly meetings with two other senior authors, until convergence where consensus was reached. Our process included the methods of literature review and reflective synthesis.
2.1 Collecting Concrete Tasks
We use the term concrete tasks to mean a prose version of a task, which may be stated using either domain-specific or abstracted language. To identify concrete tasks from previous work, the first author conducted a literature review using seed papers on network task taxonomies, visualization techniques, and surveys already known to the authors, from which a broader forward and backward citation search for relevant material followed. At this stage, 65 papers were identified. Task taxonomy papers included explicit abstract tasks that we could simply extract. For example, “Is the node an articulation node?” is listed verbatim in Lee et al.’s seminal taxonomy of network visualization tasks [2]. We also gleaned end-user tasks from non-taxonomy papers, either inferring them through definitions of graph-theoretic concepts and terminology, or through descriptions of the use cases for algorithms or techniques.
@@ 2.2 Developing Data Abstraction
As we considered which concrete tasks to include and which to exclude, we developed a data abstraction model through reflective synthesis that could serve as a filter to exclude tasks that cannot be undertaken with unlabelled graphs. It assigns graphs into four nested layers: Augmented, Attributed, Labelled, and Unlabelled, as shown in Fig. 2.
2.3 Filtering by Data Abstraction
We used our data abstraction model to filter the set of collected concrete tasks. We excluded as irrelevant any task that required information from the Augmented, Attributed, or Labelled data layers; we call such tasks labelled graph tasks. We only kept tasks that pertained to the Unlabelled layer, which we call unlabelled graph tasks.
This filtering based on our data abstraction allowed us to exclude all tasks from 24 of the papers. We chose to include at least one task from each of the other 41 papers (15 taxonomies, and 26 non-taxonomy papers). In most cases, these papers had a mix of labelled and unlabelled tasks, and we undertook fine-grained analysis to distinguish between them. As with developing the data abstraction, our reflective synthesis process involved analysis by the first author, discussed in detail with the other senior authors in weekly meetings, and iterative refinement until a consensus was reached between all authors.
For example, the Pandey et al. survey on tasks for hierarchical visual encodings of tree topology [12] contained many tasks that we excluded through data abstraction filtering by noting that they require attributes (e.g. semantic node label A, B, or C) and hierarchical wayfinding (e.g. hierarchical level Y, X, or Z from root node). To illustrate, the concrete task “is the tree balanced or un-balanced?” requires an established root node, which is a semantic label associated with a specific node, and therefore is not an unlabelled graph task. However, some tasks that appear in this survey were included by converting them into more abstract forms: for example, we converted the concrete task “count the total number of nodes in the tree” to its generic form “count the total number of nodes”. This task is equivalent to one present in Lee et al.’s taxonomy [2].
Furthermore, we used our data abstraction to glean tasks from papers that are not taxonomy papers, for example Newman’s landmark 2003 paper on complex network structure [6] implicitly suggests that “finding the length of the longest geodesic” and “counting the number of longest geodesics” are relevant tasks to draw insights about the network’s maximum traversable distances, and the induced paths that support these traversals. Both of these tasks are unlabelled graph tasks, because the length of the graph’s diameter and the number of longest shortest paths are targets that can be found without the need of semantically meaningful or persistent node labels.
Moreover, our data abstraction allowed us to proactively rule out many potential sources of concrete tasks without explicitly collecting such tasks into our lists. For example, hybrid network layouts such as NodeTrix [67] are designed for grouping nodes according to attribute values and identifying specific nodes based on semantic labels; thus, they require data at the Attributed and Labelled data layers and were not included.
Fig. 1 summarizes our results, coded by whether tasks were chosen (15), gleaned (26), or excluded (24). We also provide a spreadsheet version in Supplemental materials.
@@ 2.4 Mapping Task Abstraction
We developed the task abstraction (Sec. 5) that covered the full suite of relevant concrete tasks with a concise set of terms through multiple rounds of reflective synthesis. We finalized the framework of three levels (Scope, Action, Target) early on, but refined the set of items at each level extensively. As with previous process stages, our reflective synthesis process involved ideation from the first author, discussed in detail with the other senior authors in weekly meetings. Iterative refinement proceeded until a consensus was reached between all authors.
@@ 2.5 Analyzing Visual Encodings
We considered the strengths and weaknesses of six different visual encodings for specific concrete tasks using reflective synthesis, presented in Sec. 7. We considered the encoding results across the set of 81 eclectic benchmark networks that we collected in previous work [3]. We drew on the experience and expertise of the authors: two are doctoral students, and two are senior visualization researchers. The first author created 9-by-9 faceted panels featuring the 81 benchmark networks, one panel for each of the six visual encodings. The authors convened over weekly meetings to discuss task performance in relation to these visual encodings. These six panels are shown in Supplemental Materials.
@@ 2.6 Generating Concrete Tasks
We generated new concrete tasks for abstraction triplets where we had not found compelling examples articulated in prior work, discussed in Sec. 6. We again used reflective synthesis, spearheaded by the first author generating new tasks and the two senior authors discussing their applicability in weekly meetings.
@@ 3 Data Abstraction
Figure 2: Data abstraction model with an illustrative example. (A) In our data abstraction model, nested layers progressively decrease in semantic richness, shown here from high on the left to low on the right: Augmented is a stand-in term for named graph models with rich associated semantics, in Attributed non-topological semantic information is allowed to be associated with nodes or edges, in Labelled there are persistent unique identifiers that can convey semantic meaning, and Unlabelled is the innermost layer where all semantic information is stripped away, leaving only unique (non-isomorphic) unlabelled graphs. (B) As an illustrative explainer, drawn in red are the 6 unlabelled graphs of 4 nodes. If persistent labels (here shown as node colors) exist, as in the Labelled layer, this set of 6 graphs expands to the set of 38 labelled 4-node graphs. Within each of the 6 groupings based on graph topology (red enclosures) the graphs are structurally equivalent or isomorphic, and distinguishable from each other only through the meaning conveyed through the node labels.
3.1 Model Layers
@@
Augmented. This outermost layer of our data abstraction (Fig. 2-A) addresses the rich semantic complexity of real-world networked data; a stand-in for named graph models, which are formalisms with specific facets of associated semantic data that help us interpret and analyze the model. To illustrate, dynamic graphs have node and edge attributes parsed as time slices [43, 11, 48], compound graphs can have disjoint or overlapping groups assigned to nodes [45, 46], hypergraphs require a generalized definition of relationship that allows more than two nodes to be connected by the same edge at once [13], and heterogeneous graphs have multiple different kinds of attributes associated to nodes and edges [10, 49, 63].
@@
Augmented graphs cannot be unlabelled graphs by definition: they explicitly associate semantic information beyond purely topological structure. Any task that pertains to the use of such formalism semantics can easily be ruled out as impossible to achieve with an unlabelled graph.
@@
Attributed. Even graphs without the semantic complexity of formalisms can contain data attributes assigned to nodes and edges, which encode domain-specific information that augments the topological structure. For example, in a social network, semantic data can include node attributes such as a person’s height and age [6], while in a biological system such semantic data may take the form of edge attributes representing the reactivity of different metabolites [47]. Attribute data is used extensively in graph drawing [22].
@@
These attributes on nodes or edges are independent from and orthogonal to the topological structure of the network. Signed values (positive or negative) and scalar weights (for nodes and/or edges) are special cases of attributes. Any task that relies on the use of attributes can be ruled out as impossible to achieve with an unlabelled graph.
@@
Labelled. The next lowest layer of semantic richness is graphs with labels, which we define as semantically meaningful unique identifiers that are persistent and can serve to distinguish nodes from each other. Although this particular distinction is not common, we propose it because even graphs that are typically considered as being attribute-free in the visualization literature do in fact have labels.
@@
For instance, in the context of Graph Neural Networks (GNNs), the term ”non-attributed” graphs refers to networked data that lacks data at the Attributed layer, but does have node identifiers [68], which is data at the Labelled layer. Similarly, directed graphs require labels to establish persistent edge direction, which conversely implies that unlabelled graphs cannot be directed. By extension, unlabelled graphs cannot have self-loops.
Although labels could be considered as a special case of attributes, we distinguish them from the more general attributes in the previous Attributed layer because the Labelled layer is at the heart of the distinction between labelled and unlabelled graphs. Any task that relies on given labels is inapplicable to unlabelled graphs.
@@
Unlabelled. At the innermost layer, outlined in red in Fig. 2-A, is the connectivity information of networked data, which establishes how nodes are connected via edges, or the graph’s skeletal topology. In this layer, nodes can be distinguished from each other only in terms of topological connectivity, not through persistent names, so there is no semantic information at all at this layer. All graphs at the Unlabelled layer are topologically unique (non-isomorphic) to each other.
A specific example of a concrete task that cannot be undertaken in the Unlabelled context is find the shortest path between node A and node B. The labels of node A and node B are precisely the information not available. By extension, most pathfinding tasks are also impossible to execute in the Unlabelled context, because they require identifying specific nodes as the start or end of the induced path.
3.2 Implications
@@
To illustrate the difference between Labelled and Unlabelled layers, Fig. 2-B compares labelled and unlabelled graphs: with 4 nodes there are only 6 different unlabelled graphs (node-link views in red), while there are 38 possible labelled graphs, which we show with node colors to indicate the labels, rather than the more typical approach of using text strings (which would be too small to read). The correspondence with the unlabelled graphs is shown through the 6 groupings (red enclosures) in Fig. 2-B, each of which has a respective unlabelled graph. Although the topological structure is the same within the groups, the labelled graphs are distinguishable from each other through the persistent labels that can convey meaning. For example, in the third red enclosure from the right, with 4 labelled graphs, the single degree-3 node can be bound to each of the different labelled cases, shown with the different colors. Without labels there is no difference between the four graphs within that group, so we have only a single unlabelled graph corresponding to each grouping. In more mathematical terms, the unlabelled graph is the isomorphism class of the labelled graph [55].
@@
Within the Unlabelled layer, node identifiers are unique but non-persistent and arbitrary. However, to instantiate a graph in any computational context, such as to load networked data in memory or perform a hop-based traversal of it, there must be some kind of unique identifiers that would serve as labels during the analysis run time. To clarify, what we emphasize in this paper with our data abstraction model is not the presence of labels themselves, but whether the labels carry semantic meaning.
4 Related Work
We discuss previous work on general task taxonomies, network task taxonomies that pertain to the Unlabelled layer of our data abstraction, network task taxonomies that cover semantically richer data abstractions, and tasks from the network visualization technique literature.
4.1 General Task Taxonomies
An early taxonomy from Shneiderman [69] provides a high-level breakdown of tasks and of data types, while Amar et al. [70] provide a low-level taxonomy of analytic tasks. In 2013 Brehmer and Munzner [1] proposed bridging from low-level tasks to higher-level tasks with a multi-level typology of abstract tasks. Munzner [9] breaks down tasks into action-target pairs; we extend this approach in our taxonomy.
From this datatype-agnostic foundation, many other taxonomies focus on narrower, more specialized scopes of specific data types. For example, Valiati et al. [71] cover multidimensional data. Nusrat and Kobourov cover the intersection of geographic data and the visual encoding technique of cartograms [72]. None of these papers focus on the data type of networks, which is the subject of our work.
4.2 Network Task Taxonomies
The seminal network task taxonomy from Lee et al. [2] includes some topology-based tasks that are well-suited for unlabelled graphs, which we incorporate into our list of applicable concrete tasks. However, the majority of their topology tasks begin with a “given a set of nodes” as the starting point, which requires meaningful semantic labels that are not available with unlabelled graphs. Moreover, they do not cover the Multiple or Pair scopes at all.
Although the taxonomy by Kerracher et al. [11] primarily focuses on temporal graphs, they also discuss the more general issue that structural tasks with no attribute metadata involved are an important category, many of which we include in our list. They identify comparison as a task of interest, corresponding to our Compare action at the Pair scope. However, they do not address the Multiple scope.
Some structural comparison tasks relevant to unlabelled graphs are described by Bae et al. [14], but the main focus of their paper is the perceptual mechanisms involved in network pattern detection.
While some aspects of these taxonomies cover tasks that pertain to the Unlabelled data abstraction layer (Fig. 2-A), their main focus is tasks that primarily rely on semantic data at higher levels of semantic richness, and none have comprehensive coverage of the full set of tasks that apply to unlabelled graphs.
4.3 Labelled Graph Task Taxonomies
Many specialized task taxonomies address the types of graphs that we assign to the formalism layer of our data abstraction, including dynamic graphs that focus on changes over time of Attributed networks [43, 48, 11], heterogeneous graphs with diverse multivariate attributes [10, 49], and compound graphs with group semantics as node attributes [45, 46]. Others focus on domain-specific analysis, such as biological pathway networks [47], with tasks that rely on rich semantic attributes. However, none of these tasks are applicable to unlabelled graphs. A comprehensive 2023 survey of network visualization task taxonomies [13] illustrates how much of the previous work is dedicated to networks with attributes, confirming the need for a task taxonomy for the unlabelled graph context.
We can also explore how real-world constraints shape the kinds of tasks researchers prioritize in practice. A practitioner survey by Sahu et al. [29] highlights that large-scale network analysis often involves graphs exceeding a billion edges, with dynamic structures and complex label attributes.
4.4 Network Visualization Techniques
Visualization technique papers often include explicit lists of tasks, despite not being framed as task taxonomies. We considered many such papers, and gleaned tasks from them that appeared to fit within the Unlabelled context. For example, the GraphPrism [15] system is explicitly aimed at producing invariant plots, so its tasks were a good fit with our work. Some papers that are not aimed at invariant network visualization in particular nevertheless did serve as sources for concrete tasks, such as the Semantic Substrates [20] technique. Many technique papers list only tasks that require additional data layers, so we did not include them. For example, systems focused on the application domain of social network analysis, such as MatLink [67], rely on persistent semantic labels.